Benzene and chloroform are toxic solvents that are soluble in other. in a solution of benzene and chloroform, the mole fraction of benzene 0.450. what is the mass percentage of benzene in the mixture?
1
Expert's answer
2019-10-14T14:30:28-0400

let be:

n(mixture)=1mol, then n(benzene)=0.45mol, n(chloroform)=0.55 mol

m(benzene)=M(benzene)*n(benzene)=78.114g/mol * 0.45mol=35.151g

m(chloroform)=M(chloroform)*n(chloroform)=119.37g/mol * 0.55mol=65.653g

m(mixture)=m(benzene)+m(chloroform)=100.805g

w(benzene)=m(benzene)/m(mixture)=0.349

w(chloroform)=1-w(benzene)=0.651